### v2

# Predictive maintenance solution for industrial systems - an unsupervised approach based on log periodic power law

Abstract
A new unsupervised predictive maintenance analysis method based on the renormalization group approach used to discover critical behavior in complex systems has been proposed. The algorithm analyzes univariate time series and detects critical points based on a newly proposed theorem that identifies critical points using a Log Periodic Power Law function fits. Application of a new algorithm for predictive maintenance analysis of industrial data collected from reciprocating compressor systems is presented. Based on the knowledge of the dynamics of the analyzed compressor system, the proposed algorithm predicts valve and piston rod seal failures well in advance.
Failure Prediction; Predictive Maintenance; Time Series; Unsupervised Analysis; Renormalization group; Critical Systems; Log Periodic Power Law; Reciprocating compressors
I Introduction
Detecting the symptoms of a failure and predicting when it will occur, in multivariate or univariate time series collected via the Internet of Things (IOT), is central to the concept of predictive maintenance (PM), which is now used in almost every area of industry. PM allows a company to better prepare for a potential failure by redesigning the production process in advance or creating a workaround when shutdown is not possible. Thus minimizing the costs and the effort of standard maintenance operations through predictive engineering.
Predicting failures, provided by PM can be very profitable for a company, under the condition that PM minimizes the number of false warnings (false positives) and maximizes the number of correctly predicted events (true positives). Creating a properly working PM process faces two main problems:
  1. 1.
    related to the definition of what is a failure in the considered technological process or in the IoT data,
  2. 2.
    the development or the application of the best algorithm (based on physical description, machine learning, or statistical methods) to the data being analyzed.
While the theoretical definition of failure is well known, in practice problems are encountered with its implementation. For various reasons (economic, productive, etc.) not every failure requires corrective action. Sometimes, a minor failure is not a good reason to stop a monitored machinery or a production process. This imprecise description of failure leads to substantial difficulties when attempting to use supervised methods to build a correct Predictive Maintenance process. This would suggest that unsupervised approaches may prove to be a better suited tool for building PM processes.
This article describes an unsupervised failure prediction method patent0 used to monitor reciprocating compressor systems based on a concept used, among others, in finance, called Log-Periodic Power Law (LPPL) proposed by the Johansen2000 and Sornette2003ab . However, due to the different nature of the data analyzed, the LPPL method cannot be applied in the same way as it was introduced for bubble (or anti-bubble) detection in economic time series. Due to the different definitions of PM failure in industrial applications, modifications need to be made to the LPPL method. In the case of data describing the financial market, the variable that directly characterizes changes in the system under analysis is examined. Such a variable is, for instance the index of the analyzed financial market, or directly the price of shares.
In the case of a machine (e.g. compressor or other systems containing a number of subsystems), the analysis uses indirect data collected by sensors. It is not possible to measure the direct changes causing the failure (e.g. material degradation, cracks etc.) but it is possible to measure changes resulting from the influence of deteriorating machine components on the measured values.
In the description of the application of the presented method, which will be discussed in detail on the basis of the monitoring data from reciprocating compressors, such indirect data are, for instance, changes in the opening angle of suction or discharge valves inside the compression chamber as a function of the volume in the cylinder chamber expressed by the angle of rotation of the crankshaft. In other words, degradative (unmonitored) changes affect measured variable changes in a less visible (more distorted) ways than in the case of direct measurement of variables.
Every failure has a cause, an initiating event, let's call it an initial breakdown (IB). Using the IB concept, the failure analysis can be as follows:
  • until the IB occurs, the behavior of the machine is normal and shows no signs of failure.
  • After the IB occurs, noticeable changes begin to occur indicating future problems.
The idea of detecting the IB point in order to use it to predict failure is similar to the concept of detecting a point of trend change in a time series. Therefore, it is not necessary to predict the time of failure in the future. It is sufficient to determine whether it can be determined whether a given point in the time series is the initial (initiating) moment of IB or not. If it is known that the current point in the analyzed time series is an IB point, then, based on the knowledge of the dynamics of the monitored device, it is possible to identify the time window in the future (in units of operating time of the device) in which the failure will occur. It is impossible however, to predict the exact time of failure of the monitored device using the presented method. This is mainly due to unpredictable occurrences, such as: changes in the load of the device, periods of time when the device is switched off, etc.
The organization of the paper is as follows. Chapter II briefly describes the current status of unsupervised predictive maintenance in the industry. The chapter III answers the question of why the logarithmic-periodic oscillation detection method is suitable for detecting sudden failures in an industrial system. Chapter IV describes a numerical method for determining failure time points from the data. The part V introduces the data used for the analysis presented and the following section VI shows the results obtained using the new method. Finally, the part VII discusses the results, advantages and disadvantages of the proposed method.
II Related works
Despite the appearance of clarity, the description of a method as unsupervised is sometimes used too hastily. As described in the section (I), it is not always possible to identify with sufficient precision the time of failure as well as its cause. Even after repairs, the process of determining the time of failure or its origin is fraught with difficulties. Therefore, before describing the status of work on unsupervised methods, it should be stated, that in this work the “unsupervised method” is understood as a solution that does not require the knowledge of labelled data of any kind as well as the lack of information concerning the input data's condition as either good/healthy or anomalous.
Following the work of survey_1 (which also includes references to other research works, as well as studies of specific solutions involving a broader understanding of the "unsupervised method"), the spectrum of available solutions can be generally divided into 2 categories.
  1. 1.
    Solutions involving the use of prediction techniques to predict points in the future and then compare them with actual data to detect anomalies, as shown in method_1 . Depending on the predictive solutions used, this category of methods requires a large amount of data. In this case, a separate problem is the accuracy of the prediction part, which is a supervised method. Hence the requirement to control it, which makes the whole solution complicated if one wants to use it in practical applications.
  2. 2.
    Solutions based on measurements of distance or similarity used to evaluate the degree of data anomaly. In this case, methods using clustering algorithms method_2 or determining similarity between time series or data are used. This approach sometimes leads to problems when encountering new data the like of which did not exist in the past. At times this renders them impossible to use for real-time data analysis.
The solution proposed in this paper is an attempt to solve the issues identified in both of the above categories: the problem of data demand and the problem associated with data behavior that is not present in the past.
III The occurrence of log-periodic oscillations as a prelude to failure.
The purpose of the PM method is to provide predictions about future failures of the system described as a set of various components cooperating together. This section illustrates why the LPPL-based algorithm is applicable to failure prediction and describes the basics of the LPPL-based method.
The generalized relation describing the hazard rate $h\left(t\right)$ (or hazard function) of a certain physical quantity at time $t$ prior to material destruction by degeneration Voight1988 is of the form
$\dot{h}\left(t\right)=Gh\left(t\right)^{\delta}$ (1)
where the $\dot{h}\left(t\right)$ denotes the derivative of the function $h\left(t\right)$ in time $t$. The $\delta$ and $G$ are the parameters of model. The $G$ is a positive constant parameter. In the following, it is assumed that $h\left(t\right)$ corresponds to changes in the variable that is tracked and from which the failure is attempted to be predicted. The hazard function, based on conditional probability theory, measures the probability that the relevant variable will show signs of failure, given that the failure has not occurred before prior to time $t$.
The equation (1) has 3 classes of solutions depending on the value of $\delta$. For $\delta=1$ the exponential function is obtained
$h\left(t\right)=h\left(t_{0}\right)e^{G\left(t-t_{0}\right)}$ (2)
where $t_{0}$ is a value of the initial time.
For $\delta<1$:
$h\left(t\right)=\left[\left(1-\delta\right)\left(t-t_{0}\right)G\right]^{\frac{1}{1-\delta}}$ (3)
and for $\delta>1$:
$h\left(t\right)=\left[\left(\delta-1\right)\left(t_{c}-t\right)G\right]^{\frac{1}{1-\delta}}$ (4)
with $t_{c}$ as a constant corresponding to time in a future. As can be seen, the solutions found for $\delta=1$ (2) and for $\delta<1$ (3) do not converge for times $t>0$. Therefore, the time of failure cannot be determined from their behavior. The most interesting case is $\delta>1$ (4), where the solution has a converging point at a finite time $t_{c}$ in future. In our analysis, the time $t_{c}$ will denote a time of a potential failure and can be determined by the choice of $G=\left(\frac{1}{t_{c}}\right)^{\frac{1}{1-\delta}}$. Then $t_{c}=\frac{1}{\left(\delta-1\right)h\left(t_{0}\right)^{\delta-1}}$.
Let's assume that the degradation of a working part, whose deterioration (failure) is predicted, can be treated as a discontinuous stochastic process associated with a given monitored variable. To simplify the analysis, the second assumption is to treat the degradation changes of physical quantity $p\left(t\right)$ over time $t$ as a non-homogeneous Poisson process in which the changes occur according to the hazard rate function $h\left(t\right)$. The dynamics of such a process can be described by the equation
$dp\left(t\right)=-p\left(t\right)h\left(t\right)dt$ (5)
the solution of which can be written as
$\log{\left[\frac{p\left(t\right)}{p\left(t\right)}\right]}=-\int_{t_{0}}^{t}h\left(u\right)du=P\left(t\right).$ (6)
In our case, $P\left(t\right)$ has an approximate form
$P\left(t\right)\approx\frac{h_{0}}{\eta+1}\left(t_{c}-t\right)^{\eta+1}$ (7)
where $\eta=\frac{1}{1-\delta}$ and the $P\left(t\right)$ is shifted by the integration constant $\frac{h_{0}}{\eta+1}\left(t_{c}-t_{0}\right)^{\eta+1}$. The result obtained (7) coincides with work Ledoit2000 (compare with equation (3) in the reference 7)
Solution (7) is invariant under continuous scale invariance (CSI), which manifests itself through the scaling property of the solution $P\left(t\right)$ if the argument of the function $P\left(t\right)$ is scaled to ($t_{c}-t$). Rescaling by some factor $\nu$ the argument $t_{c}-t\rightarrow\left(t_{c}-t\right)\times\nu$ changes the solution $P\left(t\right)$ to the form $P\left(t\right)\times\mu$ where $\mu=\nu^{-1-\eta}$ IdeSornette2002 . The CSI feature, around the critical points $t=t_{c}$, is common to systems demonstrating a continuous phase transition (second order phase transition).
The basic assumption of the LPPL method is that, the described process is near the critical point of the second-order phase transition. In our case, this is the point in time at which the failure of the described system occurs. With this assumption, the final equation is obtained, the use of which for fitting is known in literature as the LPPL method Feigenbaum1996 , Sornette1996 .
Let $W\left(t\right)=\log\left(p\left(t\right)\right)$ and refer $p\left(t\right)$ to the variable by which our industrial system is analyzed as a function of time $t$. Let $t_{c}$ be the time of event that defines our phase transition in physical framework (critical point of time). Then the argument $x$ and the real function $F\left(x\right)$ are defined as
$x=t_{c}-t\textrm{ with }t<t_{c}\textrm{ and }F\left(x\right)=W\left(t_{c}\right)-W\left(t\right)$ (8)
As a starting point for our derivation, the CSI is assumed to exist around critical points.
This allows us to use the renormalization group approach, which permits us to write a certain real function $F\left(x\right)$ around a critical point $x\approx 0$ through the rescaled argument $x$ expressed by a scaling function $\phi\left(x\right)$ in the form
$F\left(x\right)=\frac{1}{\mu}F\left[\phi\left(x\right)\right]$ (9)
where $\mu$ is a constant and its argument $x$ is invariant under arbitrary linear transformation of the $x$:
$\phi\left(x\right)=\nu x\textrm{ and }x>0$ (10)
with $\nu$ as a constant.
The solution of equation (9) is the function
$F\left(x\right)=Cx^{\alpha}$ (11)
where $C$ and $\alpha$ are constants to be determined.
The request of scaling invariance (9) with the general form of the solution postulated by eq. (11) can be rewritten as
$Cx^{\alpha}=C\frac{\nu^{\alpha}}{\mu}x^{\alpha}$ (12)
which, after taking into account the identity,
$1=e^{i2\pi n}\mbox{ with }n\in\mathbb{N},$ (13)
leads to the equality
$e^{i2\pi n}=\frac{\nu^{\alpha}}{\mu}$ (14)
what allows to calculate the $\alpha$ exponent in a most general form as
$\alpha=\frac{\log\left(\mu\right)}{\log\left(\nu\right)}+i\frac{2\pi}{\log\left(\nu\right)}n.$ (15)
Therefore, the solution of equation (9) can be expressed as Nauenberg1975
$F\left(x\right)=\frac{C}{\mu}\left(\nu x\right)^{\frac{\log\left(\mu\right)}{\log\left(\nu\right)}+i\frac{2\pi}{\log\left(\nu\right)}n}=\frac{C}{\mu}\left(\nu x\right)^{\frac{\log\left(\mu\right)}{\log\left(\nu\right)}}\Pi\left(\frac{\log\left(x\right)}{\log\left(\nu\right)}\right)$ (16)
where $\Pi\left(\cdot\right)$ is a periodic function with period 11, i.e. $\Pi\left(y\right)=\Pi\left(y+1\right)$ . The index $n$ should be treated as one of the parameters characterizing the described physical system. Since $n$ and other parameters appearing in the function (16) are unknown, it is necessary to reformulate the function (16) in such a way that it can be used to fit existing data and thus determine whether a given point $t_{c}$ is a critical point.
An additional necessary condition that must be satisfied by function (16), or more precisely by its real part, is the trend that is determined by the power law ($n=0$), which is the leading order term and the oscillations associated with $n\neq 0$ will contribute as a next-to-leading order corrections.
The periodic function $\Pi\left(\cdot\right)$ can be expressed by means of the Fourier series with respect to the variable $y$ with period $T$
$\Pi\left(y\right)=\exp\left[i2\pi n\left(\frac{y}{T}\right)\right]=\sum_{k=-\infty}^{+\infty}c_{k}e^{i2\pi k\left(\frac{y}{T}\right)}$ (17)
with
$c_{k}=\frac{1}{T}\int_{-\frac{T}{2}}^{\frac{T}{2}}\left\{\exp\left[i\frac{2\pi}{T}yn\right]\right\}e^{-i\frac{2\pi}{P}ky}dy=\frac{1}{\pi}\frac{\sin\left[\pi\left(k-n\right)\right]}{k-n}\mbox{ for }n,k\in\mathbb{N}.$ (18)
However, non-zero coefficients $c_{k}$ of the Fourier expansion (17) are obtained only for non-integer differences $\left(k-n\right)$, which contradicts our claim for the existence of non-zero expressions associated with $n\in\mathbb{N}$. The problem of zero coefficients of the fourier expansion of (17) can be solved by rewriting the identity (13) to the form
$1=e^{i2\pi n}\rightarrow 1=e^{i\frac{2\pi}{q}\left(nq\right)}$ (19)
where, the variable $q$ denotes an parameter associated with the physical degradation mechanism of the described system.
In that formulation, $\alpha$ (15) can be rewritten as
$\alpha=\frac{\log\left(\mu\right)}{\log\left(\nu\right)}+i\frac{2\pi}{\frac{\log\left(\nu\right)}{q}}\frac{n}{q}$ (20)
what allows us to rewrite the equation (16) to the form
$F\left(x\right)=\frac{C}{\mu}\left(\nu x\right)^{\frac{\log\left(\mu\right)}{\log\left(\nu\right)}+i\frac{2\pi}{\frac{\log\left(\nu\right)}{q}}\frac{n}{q}}$ (21)
In this case, the expansion of $F\left(x\right)$ into Fourier series gives us the following result.
$F\left(x\right)=\frac{C}{\mu}\left(\nu x\right)^{\frac{\log\left(\mu\right)}{\log\left(\nu\right)}}e^{i\frac{2\pi}{\frac{\log\left(\nu\right)}{q}}\frac{n}{q}\log\left(x\right)}=\frac{C}{\mu}\left(\nu x\right)^{\frac{\log\left(\mu\right)}{\log\left(\nu\right)}}\sum_{k\in\mathbb{N}}c_{k}e^{i\frac{2\pi}{\frac{\log\left(\nu\right)}{q}}\log\left(x\right)k}$ (22)
where
$c_{k}=\frac{1}{T}\int_{-T/2}^{T/2}e^{i\frac{2\pi}{T}\left(\frac{n}{q}-k\right)y}dy=\frac{1}{\pi}\frac{\sin\left[\pi\left(k-\frac{n}{q}\right)\right]}{k-\frac{n}{q}}$ (23)
with a redefined variable $y=\frac{\log\left(x\right)}{\frac{\log\left(\nu\right)}{q}}$.
Given the denominator $\left(k-\frac{n}{q}\right)$ in the coefficients of $c_{k}$ (23), the main dominant terms of the series (22) are defined by the index $k$ from the set $\left\{-1+\left[n/q\right],\left[n/q\right],1+\left[n/q\right]\right\}$ where $[n/q]$ is the integer part of division $n/q$. To simplify the notation, let us redefine the index values from $\left\{-1+[n/q],[n/q],1+[n/q]\right\}$ to $\left\{-1,0,+1\right\}$.
This allows us to approximate the final form of the function $F\left(x\right)$ (22) by the first 3 largest components of the Fourier series ($c_{0}$ and $c_{-1}$ or $c_{+1}$)
$F\left(x\right)\approx{}\frac{C}{\mu}\left(\nu x\right)^{\frac{\log\left(\mu\right)}{\log\left(\nu\right)}}\left[c_{0}+c_{\pm 1}\cos\left(\frac{2\pi}{\frac{\log\left(\nu\right)}{q}}\log\left(x\right)\right)\pm\right.\left.ic_{\pm 1}\sin\left(\frac{2\pi}{\frac{\log\left(\nu\right)}{q}}\log\left(x\right)\right)\right]$ (24)
where the notation $c_{\pm 1}$ was used to denote ambiguity as to which coefficient is the second dominant one for k=1k=-1 or $k=+1$. Since only the real part of the expression is of interest (our measurements are real values), using the previous definition of the variable $F\left(x\right)$ (8) and generalizing our unknown parameters ($\mu$, $\nu$, $C$, $q$, $n$, $k$ and $W\left(t_{c}\right)$) by adding constant AA and phase $\Phi$ to the formula (24) to their new representations (AA, $B$, $m$, $C$, $\omega$, $\Phi$) one obtains the final formula which is referred to as a first-order model and used in LPPL literature Feigenbaum1996 ; Sornette1996
$\begin{split}W\left(t\right)\approx A+\left|t_{c}-t\right|^{m}\left[B+C_{1}\cos\left(\omega\log\left|t_{c}-t\right|+\Phi\right)\right].\end{split}$ (25)
For the purpose of numerical fitting the LPPL function to the data, a transformed version of the formula (25) is used, in the form of
$W\left(t\right)\approx A+\left|t_{c}-t\right|^{m}\left[B+C_{1}\cos\left(\omega\log\left|t_{c}-t\right|\right)+\right.\left.C_{2}\sin\left(\omega\log\left|t_{c}-t\right|\right)\right]$ (26)
where $C_{1}=C\cos\left(\Phi\right)$ and $C_{2}=-C\sin\left(\Phi\right)$. Both equations (25, 26) can be used to find critical time points in the time series of input data that indicate component failures in the input data.
IV Fitting method of the LPPL model to the data
[…]
Instead of trying to determine the critical time $t_{c}$ in the future, as is determined in the case of predictions of crashes in financial time series Sornette1996 ; Shu2019 ; Ledoit2000 , it is assumed that the failure happens "now". With this assumption, the calculations involved in fitting the function (26) to the data are performed for a range of time windows of different lengths from the past to "now". This corresponds to the hypothesis that the time point $t_{c}$ (time point of the phase transition) is "now" and corresponds to the last point in the input time series $t_{inp}$. Therefore, it is necessary to add to the set of parameters, an additional parameter specifying the length of the subset of time series points $t_{inp}$ preceding the time point $t_{c}$ for which the best fit of the function $W\left(t\right)$ (26) was found. The length of this subset is denoted as $l_{max}$.
In particular, our redefined set of arguments $x=t_{c}-t$ in the matching procedure is the set of points $\left<1,x_{l_{max}}\right$> in time units characteristic to the $t_{inp}$ series. The definition of the argument set $x=t_{c}-t$ in the matching procedure is replaced by the set of points $\left<1,x_{l_{max}}\right$>, where 11 corresponds to the present time and $l_{max}$ corresponds to the time of $l_{max}$ from the past. The index $l_{max}$ is the one of the parameters of the matching function (26).
As parameter fitting, the method described in the work of Shu2019 is used, appropriately modified for our purposes, i.e., by excluding the parameter corresponding to the critical time $t_{c}$ and adding the parameter $l_{max}$ to the fitting procedure.
[…]
  1. 1.
    $l_{max}$: the number of past data used for the best fit. For a small number of data there may be too many good fits (with very small fitting error), which may correspond to random correlations of the data with the form of the fitted function (26).
  2. 2.
    $A>0$: since in our case there are always positive values. It is determined by the character of the input data.
  3. 3.
    $0<m<1$: to ensure that the fitting value for the critical time $t_{c}$ was greater than zero ($m>0$) and changed faster than exponentially for times close to the critical time $t_{c}$ ($m<1$).
  4. 4.
    $2<\omega<8$: this condition avoids too fast log-period oscillations (otherwise they would fit the random component of the input data) and too slow log-period oscillations (otherwise they would contribute to the power law behaviour $\approx A+B\left|t_{c}-t\right|^{m}$)
  5. 5.
    $B$, $C_{1}$$C_{2}$: these parameters are fitted without additional constraints.
Depending on the type of input data to be analyzed, the limits of variation of the parameters to be fitted require careful adjustment.
Having determined all parameters of the fitting LPPL function, in the next step, it is necessary to determine trends based on the identified local maxima and minima of the shape of the fitted function: $T_{max}$ for local maxima and $T_{min}$ for local minima. The procedure of finding trends is carried out separately for maxima and minima in the following way:
  1. 1.
    all local extrema (N) are found,
  2. 2.
    from this set, the N-1 extreme values closest to the current time point (i.e., the point for which an attempt is made to determine whether phase transition has occurred or not) are selected,
  3. 3.
    To this set of points a straight line is fitted by linear regression. The slope of the line determines the trend for a given category of extremes (maxima or minima).
[…]
Theorem 1.
Assume that the function $f\left(x\right)$ corresponds to the best found fit of the LPPL function (26) to the analyzed input time series $ts=\{\left(t_{n}-t_{l_{max}},y_{n-l_{max}}\right),...,\left(t_{n}-t_{n-1},y_{n-1}\right)\}$ where $l_{max}>0$ is one of the fit parameters of the function $f\left(x\right)$ (26) specifying the length of the sequence preceding the actual values $\left(t_{n}-t_{n-1},y_{n-1}\right)$ defined by the index $n$. The function $f\left(x\right)$, have $N_{max}>2$ local maxima and $N_{min}>2$ local minima. Let $T_{max}$ denote the slope of the linear fit determined by the last $N_{max}$ values of the local maxima and $T_{min}$ denote the slope of the linear fit determined by the last $N_{min}$ values of local minima.
If both trends $T_{max}$ and $T_{min}$ determined for the function $f\left(x\right)$ have the same behavior (both increasing or decreasing) for the last point $\left(t_{n}-t_{n-1},y_{n-1}\right)$ of the series tsts, then the point $\left(t_{n},y_{n}\right)$ is the critical point for the series tsts. The trend of the series tsts will change to the opposite for next points $\left(t_{n+k},y_{n+k}\right)$ (where $k>0$) with respect to the trend for points preceding $t_{n}$.
In the presented method, each point satisfying fulfilling the requirements of Theorem (1) is treated as a point initiating a future failure (IB point). Determining the time of failure requires knowledge of the dynamics of the monitored system and will be discussed in the next section (VI).
V Data description
Monitoring data from a reciprocating compressor, describing the PV diagram of one of the compression chambers, was used to demonstrate the operation of the method. Based on these, the values of the angle of opening of the suction valve (OSV) expressed by the angle of rotation of the crankshaft were determined.
This value, for a given cycle described by the PV diagram, is very sensitive to changes in the amount of gas in the chamber, for example, due to its leakage through a broken valve or piston rod seal system. OSV changes are directly dictated by the thermodynamics of the compression process in the compressor chamber. Monitoring the changes in OSV provides a basis for compressor diagnostics and allows to determine compressor efficiency, valve operation, or the condition of the piston seals or piston rod sealing elements Reciprocating_compressors .
The data has been averaged to daily values and covers a period of time between 2019-08-23 and 2022-01-19.
In order to compare the results of failure prediction, data identifying the dates of repair interventions with their respective reasons for failure and the dates of observation of anomalous compressor behavior without interruption of operation were used.
VI Prediction of failures: methodology and results
To test the effectiveness of our algorithm, a backtest of the detection method was conducted on historical data starting from the initial date of 2019-10-09 ($t_{start}$) to 2022-01-01, calculating initial breakdown (IB) points for this time period. The range of the variable length of the time series $l_{max}$ was assumed to be $30<l_{max}<101$ in time units of days. Given the minimum number of observations (101 days) needed to perform calculations of the IB points, the backtest is started for time $t=t_{start}+101$ (in daily units). Then, moving forward in time to the future, the best-fit LPPL function (26) is calculated for each subsequent time $t$ by determining the goodness of fit of the LPPL function using the mean squared error (msemse).
Intuitively, one can expect that the accuracy of determining the critical points using theorem (1) will strongly depend on the error of fitting the LPPL curve (26) to the data. The smaller the fitting error, the greater the confidence that a given point of the input time series determined as an IB point according to the theorem (1) is really the IB critical point. In addition, it is expected that in the vicinity of the true IB breakpoint (before and after it), the method should find some good fits of the LPPL function with a small error, but this is not a necessary criterion for the existence of an IB point for days as time units.
The figure (1) shows 2 examples of the fit function (26) to data along with calculated criteria for trends determined from maxima and minima of the fitting function satisfying the criteria of Theorem (1).
(a) positive trends
(b) negative trends
Figure 1: Examples of LPPL function (26) fits for cases with positive trends (a) and with negative trends (b) for selected current times (dates in proleptic Gregorian ordinal units). Some fit parameters calculated for both trend categories: (a) positive trend: current date: $t_{c}=2020-08-28$, length of the time series: $l_{max}=79$, mean squared error: $mse=2.6\cdot 10^{-5}$; (b) negative trend: current date: $t_{c}=2020-09-23$, length of the time series: $l_{max}=55$, mean squared error: $mse=5.7\cdot 10^{-5}$.
The figure (2) shows the application of this procedure for the diagnosis of the critical points with additional information about the dates of failure repairs (see the description of the figure (2)).
Figure 2: Calculated IB points for the analyzed input time series (marked in the figure by a solid line with measurement points). The vertical lines of different colors indicate the diagnosed by the algorithm IB points. For each group of the IB points, the mean value of the matching error (msemse) of the LPPL function (26) is annotated. The black vertical lines indicate the dates of compressor repairs, which typically took place a few to several weeks after the algorithm detected the fault (IB points). The correlation with the diagnosed breakpoints is clearly visible, except for the prediction determined by the largest error (mse=0.000249) for 2020-12-14.
The figure (2) confirms our initial hypothesis very well. It shows:
  1. 1.
    groups of points with similar fitting errors at least 1414 days before the time of failure identification (repair is usually performed with additional delay due to the compressor operating conditions).
  2. 2.
    Dependence of the matching error on the criticality of the failure: the signaling of the grouping of critical points for the beginning of 2020-12-14 is characterized by a large error ($mse=24.9\cdot 10^{-5}$), much larger than for the group of points for which the failure has been confirmed ($mse<=5.7\cdot 10^{-5}$).
Taking these conclusions into account and comparing the recorded failures and behaviors suggesting problems in compressor operation, classified by experts as insignificant, with the predictions made by the model, it is possible to determine the threshold values of the fitting error (msemse) and the corresponding categories of predictions:
Definition 1.
Classification of calculated initial breakdown points:
  1. 1.
    Critical event: ($mse<6\cdot 10^{-5}$) severe failure expected, checking the compressor required and preparation for repair,
  2. 2.
    Monitoring event: ($6\cdot 10^{-5}<=mse<10\cdot 10^{-5}$) distinct possibility of issues, monitoring of compressor behavior required.
  3. 3.
    Irrelevant event: ($10\cdot 10^{-5}<=mse$) no significant issues predicted, compressor behaviour can be monitored though.
As shown, sometimes the algorithm detects a larger number of IB points located very close to each other. Such cases were simplified by choosing a single representation (the first IB point of the group) for each group of signals. This was done by assuming that a signal belongs to a group if its distance from the preceding signal is smaller than or equal to 33 days.
VI.1 Determining the time window of predicted failures
By comparing the failure times predicted by the algorithm with their actual occurrence, a criterion for predicting the time window in which the failure will occur can be also determined. One of the parameters for fitting the LPPL function (26) to the data is the length of the chosen sequence of data preceding the analyzed time point $l_{max}$. For the data analyzed, the time window for the occurrence of a predicted failure was defined as:
Definition 2.
The predicted time period of failure occurrence is defined as the interval $\left<n+\frac{l_{max}}{2},n+90\right$>, where $n$ and $l_{max}$ are the indices of the actual time point $x_{n}$ and the parameter defining the length of the input time series $t_{inp}$ used to find the best fit of the LPPL function (26) to a given value of $t_{inp}\left(x_{n},y_{n}\right)$, respectively.
The definition (2) is based on knowledge of the dynamics of the device for which the algorithm parameters have been defined. For other devices, all parameters should be selected based on the dynamics of their behavior. Duration of the time window with predicted failure is assumed to be valid for a certain period of time and is up to 90 days.
In the case of compressors, due to the different criticality of failures, some of them may be accepted for a longer period of time (even several months in the case of valve failures) to wait for a convenient moment of repair.
Considering:
  • the classification of alerts specified in the definition (1),
  • the selection of the representative of the groups of warnings (by selection of the initial signal for the common group of calculated IB points),
  • the definition (2) specifying the expected time window in which the failure will occur,
the raw results shown in figure (2) can be redrawn to a new form, as shown in figure (3). The correlation between predicted failure times and actual repair times, and the time periods when experts detected abnormal compressor behavior is very good for the Critical event and Monitoring event categories. Predictions in the Irrelevant event category were not confirmed by any repair and diagnosis records.
Figure 3: Redrawn prediction of failures by the algorithm in comparison with reparation times and problems detected by experts. The figure shows representations of groups of discovered IB points (colored and continuous single vertical lines) and predicted failure time periods corresponding with them according to definition (2) (corresponding colored areas bounded by vertical dashed lines). The categorization of the criticality of the forecasted problems (definition (1)) is represented by the splitting into 3 separate figures, each for a separate criticality category. The Monitoring event class remains without any marked vertical lines. The algorithm found no prediction for this category. The reparation (maintenance) dates are indicated in the figures by black vertical lines. Time periods when abnormal compressor behavior requiring monitoring and was not qualified for repair by experts are indicated by the gray colour without additional delimiting lines. The intersection of the areas determined by the algorithm as the period of failure occurrence with periods of abnormal behavior of the compressor can be seen for periods starting from 2021-07-15 in the Critical event category. The raw data of the input time series $t_{inp}$ are indicated by the solid blue line with the measurement points ($o$).
VI.2 Root cause of predicted failures
In the analysis presented here, the input data monitors the change in the opening angle of the suction valves in the compression chamber expressed by the angle of rotation of the crankshaft. The trend identified from the determination of the IB points can be used to guess the type of future failure in the compression chamber Reciprocating_compressors . By predicting the trend for times after the IB point, it is possible to try to determine approximately which part will fail - the valve or the piston rod sealing rings. Thus, when the predicted trend of the suction valve opening angle is decreasing, it is likely that the suction valve or piston rod seal rings are failing. If the trend suggests an increase in angle, this behavior indicates a leak in the discharge valve.
Thus, based on the theorem (1) and a physical interpretation based on the behavior of the time series, the algorithm is able to predict not only the time window of failure, but also the group of parts that may fail. This provides an opportunity to verify the prediction not only on the basis of event times, but also on the basis of identifying the parts that can fail.
To better illustrate the additional information regarding the location of the future failure, the figure (4) shows the same data as in the Figure (3) with additional information about the parts that actually failed, the parts in which experts have observed problems and prognosis of failures predicted by the algorithm. Details are given in the description of the figure (4).
For the entire time period analyzed, 2 cases that deviate from the diagnosis are visible
  1. 1.
    in the category Monitoring event, for date 202107152021-07-15, there is a disagreement between the predicted failure type suction valve or sealing - leakage and the diagnosed one indiation of discharge valve leakage.
  2. 2.
    the perturbation identified by experts, started in 202111302021-11-30 and identified as indiation of discharge valve leakage was not predicted by the algorithm at all.
Figure 4: The same graphs as in the figure (3) with added annotations describing the failures predicted by the algorithm - colored texts (different from gray and black), reasons for compressor reparations - black text and diagnosed abnormal compressor behavior requiring monitoring - gray text. For better visualization, areas of the diagnosed abnormal behaviour of the compressor that require monitoring are displayed in the "Monitored Event" category.
Table 1: Detailed comparison of model predictions with maintenance logs entries and expert diagnoses. The "Predictions" section contains a description of the alert: "Alert date" - date of alert occurrence, "mse" - goodness of the fit, "Predicted failure time window" - the predicted time window in which the failure may occur, and "Predicted failure" - the predicted failure diagnosis. The "Maintenance logs" section contains columns: "Maintenance date" - repair date and "Diagnosis" - the reason for the failure. The "Expert diagnoses" columns contain information in turn about: the beginning date of the anomaly notice - "Start date", the end date of the anomaly - "End date" and the symptom of the anomaly - "Diagnosis". The "Label" column contains the identification of the predicted failure. The following convention is used: True Positive (TP) event, with correctly predicted compressor failures or misbehavior, False Positives (FP), an event predicted by the algorithm that turned out to be false, and False Negatives (FN), i.e. compressor failures or instabilities that were not predicted.
Predictions Maintenance logs Expert diagnoses
Label Alert date mse Predicted time window of failure Predicted failure Maintenance date Diagnosis Start date End date Diagnosis
TP 2020-03-06 $11\cdot 10^{-5}$ 2020-04-07 - 2020-06-04 SV or Sealing 2020-04-14 SV: leakage detected - - -
TP 2020-03-27 $18\cdot 10^{-5}$ 2020-04-26 - 2020-06-25 DV 2020-05-25 DV, Sealing system: leakage detected - - -
TP 2020-08-20 $26\cdot 10^{-5}$ 2020-09-23 - 2020-11-18 SV or Sealing 2020-11-10 Sealing system failure; SV and DV: leakage detected - - -
FP 2020-12-14 $249\cdot 10^{-4}$ 2021-01-19 - 2021-03-14 DV - - - - -
TP 2021-06-18 $36\cdot 10^{-5}$ 2021-07-22 - 2021-09-16 DV 2021-08-11 DV - leakage detected 2021-07-06 2021-08-02 DV leakage
FP 2021-07-15 $36\cdot 10^{-5}$ 2021-08-26 - 2021-10-13 SV or Sealing - - 2021-08-17 2021-10-21 DV leakage
FN - - - - - - 2021-11-30 2022-03-02 starting sealing leakage
VII Discussion
The comparison of the information resulting from the failure predictions and comparing it with the knowledge from maintenance logs and expert detection of periods of anomalous compressor operation, including the prediction classification, is provided in Table 1.
Major challenges in the field of industrial application of failure prediction, especially in the unsupervised version, is a number of issues relating to proper identification of failures and behaviors of monitored devices. These difficulties mainly stem from:
  • the large variety in the types of failures,
  • the small number of failures compared to the amount of data,
  • the large variety of behaviors leading to the same type of failure,
  • the difference in operating conditions, which can cause ambiguity in labeling accurate data.
The problems are very difficult to solve if the methods used to predict failures are based on the analysis of numerical values of input data (by calculating similarities, correlations, logic trees, building naural networks, etc.). Normalization and/or standardization procedures only introduce a common scale to the analyzed data.
The proposed method introduces a new type of procedure, which is based on the search for common functional behavior (26). From the point of view of the numerical values of the analyzed data, the course of the fitted function can be very different for different events - different patterns of the function for different values of fitted parameters. Even when new data appears with values that did not exist in the past, it is possible to determine the critical point (IB) for a potential event, as the model fits functions to the data.
The IB points determined by this method, upon which the time window of failure occurrence is predicted in the next step, are the trend change points in the data. From this perspective, the described solution can be reduced only to the task of determining the trend change points.
To calculate the key performance indicators (KPIs) of the presented algorithm, the generally known indicators can be used: $Precision=TP/\left(TP+FP\right)$ and $Recall=TP/\left(TP+FN\right)$, where TPTP - defines True Positive events, i.e. correctly predicted failures or abnormal behavior in the operation of the compressor, FPFP - False Positives, i.e. events predicted by the algorithm that turned out to be false and FNFN - False Negatives, i.e. failures and instabilities of the compressor that were not predicted.
In the analysis of results (section VI), the limits of acceptance of errors of fitting the LPPL function (26) determining the criticality of the predicted events (definition 1) were defined on the basis of the results. Therefore, in order to calculate the PrecisionPrecision and RacallRacall indicators, all results are taken into account without distinguishing them due to the defined criticality.
Comparing the predicted failure periods, taking into account the dates and predicted types of failures with the dates and descriptions of Maintenance logs or recorded faults (Figs. 3 and 4 and the table 1), the calculated values of TPTP, FPFP and FNFN are as follows: $TP=4$, $FP=2$, $FN=1$ . Hence $Precision=0.67$, $Recall=0.8$. Given that, this result takes into account the dating of the predicted failure period along with the prediction of the cause of failure, such an outcome is considered very good.
In summary, the list of advantages and disadvantages of the presented method is a consequence of a paradigm change in data behavior classification, from the one based on numerical values to the one based on functional similarity.
Advantages of the method:
  • the proposed model can be applied to very short time series (in our case, the minimum length of the series is only 101 points).
  • There are no problems with data that appears for the first time. In the proposed solution, the part that qualifies certain data as IB is based solely on functional behavior. This universality is due to the renormalization group approach.
  • The simplicity of the final production solution. The most difficult part of the algorithm is fitting the function to the data. Since the method does not contain components based on supervised methods, there is no need to monitor their quality.
Disadvantages of the method:
  • data: the method is applicable to data that describe a physical process that degrades/changes due to perturbations introduced by interacting elements. This is because the method searches for behavior characteristic of phenomena in which phase transitions can be observed. Hence, not all data are appropriate for the described method.
  • Matching the function to the data is based on the proper determination of boundaries of the parameters to be matched (26). This requires individualized adjustment of the ranges of change of these parameters to the monitored device.
  • It is required to select the time step of the input time series in such a way that it is consistent with the dynamic characteristics of the monitored device.
VIII Conclusions
This paper presents the application of a methodology for describing critical behavior in complex systems based on the renormalization group approach in unsupervised predictive maintenance. The proposed algorithm analyzes the behavior of a complex system based on a time series representing the physical behavior of the system. To demonstrate the effectiveness of the algorithm for industrial applications, predictive results are presented for time series describing the thermodynamics of the gas compression process in a monitored reciprocating compressor in one of the compression chambers.
It was shown that failures in the analyzed industrial system can be treated as critical behavior in complex systems. Then the symptoms of future failure appear in the form of Log Periodic Power Law structures for phase transitions in the analyzed time series. Based on the most generalized scheme for describing the behavior of an analyzed system in the vicinity of phase transitions of the 2nd kind based on the Log Periodic Power Law, a new way of predicting failures in compressor systems is proposed.
The presented algorithm is based on 3 steps. In the first step, the algorithm determines the IB points in the analyzed time series by means of fitting the LPPL function (equation (26)) using the proposed theorem (1). The second step of the proposed method is based on the knowledge of the dynamics of the monitored system, and specifies the time window in which the predicted failure may occur. In the last step, a criticality classification of the predicted failure is carried out, based on the goodness of fit of the LPPL curve to the data (critical event, monitoring event, insignificant event).
Taking into account the specificity of problem detection in industrial systems (the demand to reduce the number of false alarms and to minimize the number of unpredicted events), it has been demonstrated that it is possible to experimentally determine such an error threshold of fitting the LPPL function to the data that all serious failures can be predicted if the fitting error is smaller than the threshold. In addition, it is also possible to define such thresholds for the LPPL curve-fit error to the data, for which the area of occurrence of less critical failures that do not require rapid intervention can be defined.
The method can also be applied to predictive IoT analysis of other industrial systems.